nazia is twice old askanwal. if 10 years ago, nazia was 4 times as old as kanwal, then the present ages of nazia and kanwal are?
Answer by ewatrrr(24785)


Hi
nazia is twice old as kanwal
Let x and 2x represent as Kanwal and Nazia's ages NOW respectively
Question states*** 10years ago...
2x-10 = 4(x-10)
Solving for x
30 = 2x
x = 15, Kania age Now. Nazia is 30. (2*15)

CHECKING our Answer***
20 = 4*5